OpenStudy (natsent06):

NEED ANSWER FAST WILL GIVE MEDAL A blueprint has a scale of one - half inch to one foot. On the blueprint, one wall me asures six inches in height. If H represents the height of the actual wall, what is H?

OpenStudy (dbzfan836):

hmmm, I dunno

OpenStudy (dbzfan836):

@sourwing we need help here

OpenStudy (natsent06):

yea we need help

OpenStudy (dbzfan836):

@sourwing come quick

OpenStudy (anonymous):

1/2 inch (scale) = 1 foot (real) 6 inches (scale) = x feet (real) Just cross multiply and you'll have your answer (1/2)x = 6 solve for x

OpenStudy (dbzfan836):

x=6

OpenStudy (anonymous):

Don't forget to medal + fan pls

OpenStudy (dbzfan836):

Can I also heve medal

OpenStudy (anonymous):

Hell no, also your answer is wrong

OpenStudy (dbzfan836):

hey, noc foul language. your account can be banned

OpenStudy (dbzfan836):

then u lose every thing

OpenStudy (anonymous):

so scared btw divide both sides by 1/2 and it gives you x = 12 feet

OpenStudy (anonymous):

12 feet